IN right angle triangle ABC angle B is 90 degree and D is the midpoint of BC.prove that BC 2=4(AD 2-AB 2).

Grade:9

1 Answers

Abhirup Roy
7 years ago
ABC is a right angled triangle right angled at B. D is the midpoint of BC and A and D are joined.
By Pythatgoras’ theorem
AD2=AB2+BD2 or, BD2=AD2-AB2 or, (1/2BC)2=AD2-AB2 or, 1/4BC2=AD2-AB2 or, BC2=4(AD2-AB2)
                                                                                                                                           proved